How frequency distribution might be used to understand and improve inventory systems?
Frequency distribution, helps us to understand the trends of data. If gives an idea about how data is spread. To graphically represent the data, a histogram is used.
In an inventory system, this measure helps in understanding the demand for different classes of items, based on which inventory ordering can be adjusted.
example: In in a shoe store, if the frequency distribution shows that, shoe of size 9 is sold maximum then the store can order more of size 9 shoes.
How many different hands of a 13-card might you have out of standard deck of 52 playing?
You can draw C(52,13) = 52! /13! 39! = 635 013559 600 different 13-card hands from a deck of 52 cards.
What is the probability of drawing a black king from a standard deck of 52 playing cards?
2 in 52 is the chance you would draw a black king in a deck of normal playing cards.
When could you use theoretical probability?
When there is a good theoretical model for the experiment and the model allows you to identify all the factors affecting the outcome and determine their impact on the outcome. Even if you cannot identify all the factors, you can still use theoretical probability but the predictions from your model will be less reliable. Econometrics is a good example of using theoretical probability based on an incomplete understanding of the model.
What is the probability of rolling a sum of more than 4 with two number cubes?
With two six-sided dice, there are 36 possible outcomes. Let's look at the outcomes which the sum is less than or equal to 4: {1.1 1.2 1.3 2.1 2.2 3.1} That's 6 outcomes, which leaves 30 outcomes with greater than 4. So 30/36 = 5/6 or 83.333%
use binomial probability. (nCx)(px)(qn-x)...... N=trials x=outcomes p=probability in one trial. q=1-p...... N=20 x=4 p=1/6 q=5/6.......... nCx or nCr is a function on any scientific calculator, may need to search for it..
(20C4)(1/6)4(5/6)16=.2022 or 20.22%
__________________________________________________________________
The above probability is the theoretical probability. The experimental probability would be: P(3) = 4/20 = 0.20 = 20%.
What is the probability that the sum of the dice is less than 6 and one die shows a 3?
The probability that the sum of two dice rolled is less than 6 when one die comes out
3 is 2/6 = 1/3, since there are only two outcomes that the second die has so the sum turns out less than 6. these are (2) and (1), each with the probability of occurring of 1/6.
What is the chances you will get a blue marble if there is 4 marbles in all?
Depends on the color of the marbles, or more specifically, it depends upon how many of the marbles are blue. If 0 marbles are blue, then the chance is 0. If 1 marble is blue, then the chance is 1/4. If 2 marbles are blue, then the chance is 1/2. If 3 marbles are blue, then the chance is 3/4. If 4 marbles are blue, then the chance is 1. These probabilities assume that you draw only once. The other possibilities are that you draw twice, or three times, or four times. If drawing multiple times, it can be done by putting back each marble after it was drawn, or by keeping it out after it has been drawn.
What is the chance of rolling a number above 3 on a dice?
Six side dye. Let x be the random variable, x ={1, 2, 3, 4, 5, 6). The probability for any of the six possible outcomes is P(x) = 1/6. This gives a flat probability distribution.
The probability that the outcome in the rolling of one dye is 3 or a smaller number
(2 or 1) is; P(x≤3) = P(x=3) + P(x=2) +P(x=1) = 1/6 + 1/6 + 1/6 = 3/6 = 1/2
The probability of the complementary event, that the resulting event is a number above 3 is: P(x>3) = 1 - P(x≤3) = 1 - 1/2 = 1/2 = 0.50 = 50%
What is the chance you will die today?
Same as anyones: likely.
There's probable over 1,000 ways you can die in less than a hour.
-Injuries
-diseases.
-suicide.
-murder.
-heart frailer.
-and more.
It is a measure of the spread of the outcomes around the mean value.
The probability is 0.09766%.
Each toss has a ½ chance to be heads. To combine probabilities use multiply them. So the probability to get two heads out of two tosses is ½ * ½, and three heads out of three tosses is ½ * ½ * ½. So the exact answer is 0.5^10
What is the probability of picking a two replacing it in the deck of cards and then picking a king?
there is a 1 in 13 chance of picking a 2 in a deck of cards, then picking a king afterwards, its also a 1 in 13 chance, but together, theres a even slimmer chance.
top, opt, pot
What is the probability that a far six-sided die rolled until the first 5 occurs on the tenth roll?
Assuming that the die is fair rather than far, the probability is (5/6)9(1/6) = 1,953,125/60,466,176 = 0.0323 approx.
How many 10's are in a deck of 52 cards?
4
a 10 of hearts, a 10 of spades, a 10 of clubs, and a 10 of diamonds
What is the probability of rolling a number that is a multiple of 3 on a fair number cube?
A fair cube has six faces numbered 1, 2, 3, 4, 5, 6
3 and 6 are multiples of 3
so you have 2 chances in 6
or 1 in 3 in it's simplest form
What is the probability tossing 3 nickels?
Depends on how many nickels you have. If you have less than three nickels the probability is zero. If not, 100% (should you decide to throw them).
What is an advantage of stereotactic radiosurgery?
The beam can be focused on any structure in the brain, allowing access to tumors or malformed blood vessels that cannot be reached by open-skull surgery.
What is the number of elements in an event space for rolling an even number on a single die?
The even numbers that can be rolled on a single die are 2,4, and 6 so the number of elements in the event space is 3. X={2,4,6}.
What is the probability of flipping 4 fair coins 1 time and getting all tails?
The probability of getting all tails when you flip 4 fair coins 1 time is:
P(4 tails) = (1/2)4 = 1/16 = 0.0625 = 6.25%
If you were throwing one numbered cube (dice) with each face numbered 1 to 6, then the chance of getting a number greater than two is 4/6 (or 2/3) since there are 4 numbers greater than 2, and six numbers in total.
Similarly, if you were throwing one dice, the chance of getting a number less than 2 is 1/6.
To determine the probability of getting both of these outcomes on two dice (greater than 2 on one dice and less than 2 on the other) we simply multiply these two values together.
2/3 x 1/6 = 2/18 = 1/9.
So there is a one in nine chance of this outcome occuring.
What is the probability for family to born 2 boys and a girl?
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability of the other two being boys is 0.4994